### v3

# Intents, Techniques, and Components: a Unified Analysis of Interaction Authoring Tasks in Data Visualization

Abstract.
There is a growing interest in designing tools to support interactivity specification and authoring in data visualization. To develop expressive and flexible tools, we need theories and models that describe the task space of interaction authoring. Although multiple taxonomies and frameworks exist for interactive visualization, they primarily focus on how visualizations are used, not how interactivity is composed. To fill this gap, we conduct an analysis of 592 interaction units from 47 real-world visualization applications. Based on the analysis, we present a unified analysis of interaction authoring tasks across three levels of description: intents, representative techniques, and low-level implementation components. We examine our framework’s descriptive, evaluative, and generative powers for critiquing existing interactivity authoring tools and informing new tool development.
Visualization, Theory, Interaction Authoring
copyright: nonecopyright: none
© 2025 All Authors. Version 3 (v3) supersedes v1 and v2, which used different license and acknowledgment terms. Please do not use v1 or v2.
1. Introduction
As the support for creating static visual representations of data has matured in visualization grammars and authoring tools (liu_data_2018; ren_charticulator_2018; satyanarayan_lyra_2014; tsandilas_structgraphics_2020; wang_falx_2021), there is a growing interest in enabling easy interaction authoring— adding interactive behavior to static visualizations—through declarative languages and interactive systems (satyanarayan_vega-lite_2016; zong_lyra_2020; latif_kori_2022; sultanum_leveraging_2021; morth_scrollyvis_2022). To understand the expressiveness of extant interaction authoring systems and to inform the design of new languages and tools, it is important to have theories and models that describe the task space of interaction authoring in data visualization. While previous research has provided taxonomies and frameworks (yi_toward_2007; heer_interactive_2012; munzner_multi-level_2013; gotz_characterizing_2009) to describe tasks in interactive visualization, they primarily focus on how visualizations are used for analytical or communication purposes, rather than how interactivity is composed. In this paper, we address this gap by providing a unified framework that describes interaction authoring tasks at multiple levels of abstraction.
To understand what level of abstraction means in the context of interaction authoring, consider the scenario where an author wants to add interactivity to a static visualization consisting of multiple line graphs showing stock prices of various companies over a period of time. At the lowest level, the author may specify that when the user draws a path, the visualization should update to show only the line marks with trends similar to the path drawn. This concrete specification may be composed using a programming language. At a higher level, this specification can be more abstractly described as “query sketch” (wattenberg_sketching_2001), which may be provided as a standard template in an authoring tool with graphical user interfaces. At the highest level, the author may have the option to just state that she wants to add filtering support to a recommendation system, which can generate applicable techniques and corresponding low-level implementations for her to choose from.
To analyze such multifaceted interaction authoring tasks, we first identify three levels of abstraction in interaction authoring: intents, techniques, and components. We then classify 592 interaction units from 47 real-world visualization applications into categories at the intent and component levels. With the resulting intent and component taxonomies, we identify representative techniques for each intent category and required components for each technique in a unified analysis.
Through the analysis, we make the following contributions:
  • We reinterpret and consolidate existing taxonomies on user intent. Based on the refined user intents, we define four types of authoring intent: Connects visual objects and changes in visual properties, Connects visual objects and changes in visible area, Connects visual objects and changes in visual mapping, Connects visual objects and changes in data.
  • We consolidate techniques mentioned in existing taxonomies and extend the lists of techniques based on example analysis.
  • We identify a set of low-level primitives for composing interactivity, organized into three categories: representation components (e.g., visual object, channel, encoding, scale), observable interaction components (e.g., event, hit object, target), and internal components (e.g., camera, state variable, target evaluator).
  • We contribute a unified framework of the task space of interaction authoring across three levels of abstraction: intents, techniques, and components (Figure 1). For each type of user intent under an authoring intent, we identify representative interaction techniques; for each technique, we describe how the desired behaviors can be achieved through the low-level components.
We discuss the descriptive, evaluative, and generative powers of our analysis. In particular, we discuss the implications of our work on the design and evaluation of interaction authoring tools. As part of the third contribution, we build an interactive web application that enables visualization researchers to collect real-world examples collaboratively. The website can be found in our supplementary material 111supplementary materials.
2. Background
2.1. Interaction Authoring
Interaction is a frequently used term in a wide range of literature on HCI and Visualization. These two survey papers (hornbaek_what_2017; dimara_what_2019) introduce seven concepts (i.e., Dialogue, Tool Use, Optimal Behavior, Transmission, Embodiment, Experience, and Control) of interaction. Among the concepts, Dialogue (hornbaek_what_2017; dimara_what_2019) defines interaction as a cycle of communication acts channeled between human (via perception and action) and a system (via input and output). By drawing from Norman’s seven stage model (sedig_interaction_2013), this concept is emphasized with mapping that a user seeks for a way to achieve their intention with tasks that a user translates their goals to a intended system state. In addition, tools or models for constructing interactions (leiva_enact_2019; blouin_interacto_2021; martin_causette_2022; beaudouin-lafon_instrumental_2000) introduce how they define interaction and how their interaction building system works. Leiva et al. (leiva_enact_2019) presents a prototyping tool for touch-based mobile interactions. They define interaction as a set of rules to associate user inputs with system outputs. To elaborate, user inputs (i.e., touches, movements, expressions, voice) captured by devices are translated into system outputs (i.e., UI feedback, underlying operations), resulting in visual outputs through feedback from corresponding operations performed by the underlying system such as databases or servers. To construct such interactions, the tool user employs a first-class object called touch events. the user can manipulate touchable UI elements or program functions to establish interactive behaviors (i.e., touch-stamp, pinch, drag, touch), thereby establishing connections between input and output properties (i.e., input: touch, output: the position change of an item). Similarly, other systems define interactions in the same manner: to support dialogues between a system and its users who engage with a UI to control the underlying system (blouin_interacto_2021), to establish causal relationships, such as input events instigating state changes or the propagation of values (martin_causette_2022), and to describe methods to integrate techniques by delineating the feedback from interactions based on the users’ perspective (beaudouin-lafon_designing_2004). While the existing body of literature and systems depict interaction and the process of constructing it as connecting humans and systems across various stages from input to output, this concept is not limited to data visualization.
The survey paper (dimara_what_2019) argues the importance of involving Data entity and Data-related intent to the definition of interaction in visualization. For example, the entity for interactive data visualization would be data, computer, and human. It is because visualization is for knowledge construction or sense-making data rather than the easiness or intuition of using a system. In addition, visualization, as a data interface, enables human to manipulate an intangible data sources, which can be linked to data-related user intent. This data-oriented interpretation of interaction is also seen in visualization authoring tools and languages (bostock_d3_2011; satyanarayan_vega-lite_2016; lu_automatic_2021; zong_lyra_2020; satyanarayan_reactive_2015).
Upon reviewing the literature, we discovered that connecting tasks at multi-level abstraction are crucial for interpreting interaction authoring. Consequently, this paper considers Interaction Authoring in Visualization as a process of connecting a user with a system by integrating techniques, system components, and data throughout the entire interactive visualization pipeline.
2.2. Existing Frameworks and Taxonomies in Interactive Visualization
Describing human behavior at multiple levels of abstraction is a cornerstone of activity theory (brown_perspectives_1999; engestrom_activity_1999), which has been informing research and design in human-computer interaction (kuutti_activity_1996; nardi_activity_1996; nardi_context_1998). Gotz and Zhou (gotz_characterizing_2009) adopt activity theory to describe four levels of abstraction in visual analytics. Pike et al. (pike_science_2009) identify high-level and low-level elements of interaction in terms of both user tasks and visualization features. Roth (roth_empirically-derived_2013) develop a multi-tier taxonomy for cartographic interaction primitives at multiple levels: goals, objectives, operators, and operands. Brehmer and Munzner (munzner_multi-level_2013) present a multi-level typology that describes the why, how, and what dimensions of user task in visualization. These works emphasize the importance of having a holistic understanding across multiple levels of abstraction, but they all focus on tasks in using a visualization, and does not address the creation of interactive visualizations.
Taxonomies and models of interaction in visualization also exist at individual levels. Yi et al. (yi_toward_2007) introduce a taxonomy of user intent, by raising a need to connect user goals with interaction techniques. Pike et al. (pike_science_2009) also discuss interactions at the high-level (user intents) and the low-level (techniques). In addition, Heer and Shneiderman (heer_interactive_2012) present categories of interactive dynamics capturing important user tasks. The works offer an understanding of the design space of interactive visualization at the user intent or techniques level.
The early literature also introduced taxonomies at the individual level. Dix et al. (dix_starting_1998) present four classes of tasks for incorporating simple interaction into static visualization, Buja et al. (buja_interactive_1996) define three classes of view manipulation along with corresponding tasks for high dimensional data visualization, and Shneiderman (shneiderman_eyes_2003) suggests seven tasks based on data types. However, these taxonomies only cover different classifications at the technique level, necessitating the consolidation of existing taxonomies. (schulz_design_2013; rubab_examining_2021) At the implementation level, Chi et al. (chi_operator_1998) offer a taxonomy based on operators and operands. Vega-Lite (satyanarayan_vega-lite_2016) presents a declarative grammar that describes how to compose interactivity with primitives; however, they focus more on the grammars’ capabilities rather than linking different levels. Animated Vegalite (zong_animated_2022) discusses unified interaction and animation tasks based on their similarities at high and low levels of abstraction. Reactive Vegalite (satyanarayan_reactive_2015) presents a comprehensive data model demonstrating the entire data transformation process, from user input to output, in interactive data visualization. However, these two works primarily focus on lower levels, minimizing the involvement of human activities such as user intent.
Our work seeks to reinterpret and link these separate single-level accounts in a unified analysis focusing in interaction authoring tasks, rather than communication. In trying to connect these multiple levels of abstraction, we also modify some of these classifications to derive new concepts and taxonomies for interaction authoring.
3. Choice of Abstraction Levels
We decide to perform analysis on interaction authoring tasks at multiple levels of abstraction because current literature has pointed out that a single level of abstraction is insufficient to capture the rich semantics of visualization activities in terms of motivation, context, goals and operations (pike_science_2009; roth_empirically-derived_2013; munzner_multi-level_2013; gotz_characterizing_2009). The scenario in the introduction illustrates the importance of supporting interaction authoring at multiple levels of abstraction. Although a few multi-level visualization frameworks exist (e.g., Brehmer and Munzner (munzner_multi-level_2013), Gotz and Zhou (gotz_characterizing_2009)), we did not closely follow their level characterizations because of our distinct focus on interaction authoring. In this section, we define the three levels of abstraction (i.e., intent, technique, and component) in our analysis and explain why they are chosen by situating them in the context of existing literature.
Intents. In our framework, “intents“ include “authoring intent“ and “user intent“. Firstly, “authoring intent” is what an author wants to achieve when authoring interaction. We define “authoring intent“ based on “user intent“ from existing literature, and the definition of interaction authoring in the section 2. Yi et al. (yi_toward_2007) emphasize the importance of linking “user intent“ with techniques, demonstrating the impact of “user intent“ throughout the data exploration process. As a result, we regard their taxonomy as a primary reference when consolidating “user intent“. They elaborate that different interaction designs can be interpreted to achieve the same goal, called “user intent“. For example, consider the following two designs: (1) clicking on a circle in a scatter plot changes its color, and (2) swiping horizontally over a bar chart on a phone fades the bars outside the swiped region into the background. Both interaction designs serve the same intent: to select items of interest. Besides Yi et al.’s taxonomy, Heer and Shneiderman’s interactive dynamics categories (heer_interactive_2012) and Brehmer and Munzner’s how level (munzner_multi-level_2013) are also operating at this level of abstraction for “user intent“. The former presents a more extensive list of categories, while the latter not only addresses the user intent but also examines the relationships between user intents and other levels. Taking these interpretations into account, we define “authoring intent” that is to seek a task to translate these “user intents“ to desired system elements. For example, to enable users to select a bar, an author wants to connect a bar and corresponding changes in its visual properties by programming.
Techniques. In our framework, techniques are to achieve a user intent, which is ultermately for the goal of fulfulling authoring intent. Techniques often tightly coupled with the type of items being manipulated. For instance, the two designs in the previous paragraph are examples of two different techniques: point select and range select. The former is about selecting a single item, while the latter is about selecting items within a specific pixel range. For the same reasons associated with the intents, we regard the taxonomies by Yi et al. (yi_toward_2007) and Heer and Shneiderman (heer_interactive_2012) as our primary references. These taxonomies feature representative techniques for each high-level user intent. We consolidate their techniques and add more techniques discovered while analyzing examples.
Components. Finally, components are the building blocks to implement interaction designs. We consider this level similar to the what level in Brehmer and Munzner’s typology (munzner_multi-level_2013). Their account at the what level, however, is not detailed and does not address interaction authoring. In general, there is little work that systematically identifies the primitives in authoring interactive visualizations. A represented relevant work is Vega-Lite (satyanarayan_vega-lite_2016), which identifies what comprises interactivity from the perspective of declarative specification. Our analysis at this level seeks to provide more details on what inputs and outputs are involved to translate UI events into visual changes.
Given these three levels of abstraction, we provide an overview of the analysis methods and process used to identify the specific categories at each level in Section 4. Section 5-8 then present the results of the identified categories at each level and a unified account of how these categories relate to each other across levels.
4. Analysis Method and Process
The primary goal of our analysis is to identify common categories at each of the three levels and organize these categories across levels. Similar to previous work that focuses on identifying high-level intents (yi_toward_2007) and low-level primitives (liu_data_2018; thompson_understanding_2020) in data visualization, we ground this work in the analysis of real-world interactive visualization examples.
Example Collection. We collected examples based on the following considerations. First, the static visual representations should encompass a variety of chart types and designs. Second, the interactive behavior should go beyond minimal features such as showing tooltip on hover. Third, we want to sample visualizations designed for both exploratory analysis as well as storytelling.
Based on these criteria, we identified 47 real-world interactive visualization examples from 16 online sources, including web applications (e.g., GOV||DNA (sudox_govdna_nodate), ESA Star Mapper (interactive_esa_nodate)), news outlets (e.g., the New York Times (the_new_york_times_new_nodate), the Economist (the_economist_big_nodate)), digital publishers (e.g., the Pudding (the_pudding_pudding_nodate), Quartz (quartz_quartz_nodate)), design galleries of visualization toolkits (e.g., bl.ocks (popular_blocks_popular_nodate)) and systems (e.g., Tableau Public (tableau_tableau_nodate), datasketch (datasketch_datasketch_nodate)), commercial demos (e.g., heavy.ai (heavyai_1990_nodate), Gramener (gramener_gramener_nodate)), and portfolio sites of design agencies (e.g., Column Five Media (column_five_media_column_nodate)).
In terms of chart design, 15 out of 47 examples consist of multiple views, which are arranged either as panels laid out in a dashboard, or as frames in a scrollytelling format, or as tabs activated using toggle buttons. In total, the example collection represents 20 different chart types: scatter plot, symbol map, heat map, choropleth map, streamgraph, bubble chart, sunburst chart, index chart, simple/grouped/stacked bar chart, (stacked) area chart, dumbbell chart, radar chart, 3D particle map, treemap, 3D area chart, network map, and line chart.
Assumptions. Given our focus on interactivity, we assume that the static representations have been created. If a visualization component does not change in response to user action, it is not included in the analysis.
Unit of Analysis: Event & Target. Since many examples are quite complex, involving multiple views responding to multiple user actions, we chose an event-target pair as the unit of analysis: an event refers to a keyboard or a mouse event, and a target refers to a set of visual objects exhibiting visible changes as a result of the event. In Figure 2, we have two targets for the hover event: (1) all the circles, and (2) the tooltip. This example thus is decomposed into 2 units of analysis. In total, we analyzed 592 units of analysis in the 47 examples.
Figure 2. Hovering to highlight a circle mark and show a tooltip
Process Overview. We conducted the analysis in three phases. We focus on intent categories and component categories in the first two phases, and identify representative techniques in the third phase.
Phase 1: We first adopted a top-down approach to apply existing taxonomies to the examples. At the intent level, we tried to assign each unit of analysis a category by consolidating Yi et al.’s taxonomy (yi_toward_2007), Heer and Shneiderman’s taxonomy (heer_interactive_2012), and Brehmer and Munzner’s how classification (munzner_multi-level_2013). At the component level, we analyzed each unit of analysis by labeling the selection primitives in Vega-Lite’s interaction grammar (e.g., type, predicate, domain—range, init, transforms, resolve). At the early stage, All three authors independently coded the examples and met to resolve disagreements. Initially, discussions on a single complex example could take an hour due to differing interpretations of intent categories and selection primitives. However, as we aligned on the concepts, we could review three examples within an hour. After the initial ten examples, two authors continued coding and involved the third only when needed for disagreements.
Phase 2: As we made progress in the first phase, we also discovered potential issues that prevent us from directly using existing classifications for interaction authoring. Section 5-7 elaborate on these issues. We thus adopted a bottom-up approach to revise and more precisely define the categories. At the intent level, we paid close attention to ambiguous cases, and analyzed the similarities and essential differences between related examples to revise the intent categories. At the component level, we identified both components that are observable to users as well as those that are internal in a program. Section 5-7 provide a detailed account of the results from this phase, along with the rationales for choosing the intent and component categories.
Phase 3: With the revised intent categories and component types, we performed labeling adjustment. After adjusting labels at the intent level, we grouped similar units of analysis and identified common techniques associated with each intent. We also identified common components used in each technique. Section 8 provides a unified account of the results from this phase.
5. Intent Categories
In this section, we discuss the issues and findings related to analysis at the intent level.
5.1. Issues with Directly Using Existing Taxonomies for Interaction Authoring
As mentioned in Section 3, we used three existing taxonomies (yi_toward_2007; heer_interactive_2012; munzner_multi-level_2013) to label intent. While the taxonomies were highly useful in providing conceptual scaffolding in our analysis, we did encounter issues that prevent us from directly applying these existing taxonomies to interaction authoring: focus on user intent instead of authoring intent.
5.1.1. User Intent vs. Author Intent
The existing taxonomies describe intent in terms of what users want to achieve. This is evident in (yi_toward_2007), where the meaning of each intent is explained in the form of “show me …”, and in (heer_interactive_2012), where the “categories incorporate the critical tasks that enable iterative visual analysis”. However, to enable users to achieve their goals, authors need to first articulate what they want to achieve in terms of how the target visual objects should change. In other words, what to connect to change the target visual objects. For example, in Figure 2 the user intent here is elaborate using Yi et al.’s taxonomy (yi_toward_2007), which describes what users want to achieve: to get more information about the station. To support this user intent, authors need to specify how the visibility and position of the tooltip should be updated based on mouse events, as well as how the tooltip content should be updated by retrieve the corresponding data. To do so, they should intend to link the input event, the tooltip, and its visual properties. The authoring intents here are very different from the user intent elaborate.
(a) Distribution of visualization designs created by participants by design use frequency.
(b) Distribution of participants by tool type and frequency of use.
Figure 3. Interactions involving coordinated views.
5.1.2. Inconsistencies within the Same Taxonomy
We also found that the categories from the same taxonomy can focus on different aspects of interactivity. Consider the two examples in Figure 3. Figure 3(a) (liu_immens_2013) shows two bar charts, brushing over a month bar in the left chart will show the daily distribution of that month in orange in the right chart. Figure 3(b) (the_economist_big_nodate) shows a list of currencies in the left table, with a base currency (US dollar) colored in dark gray. The lollipop chart on the right shows all the currencies’ Big Mac index against the base currency. Selecting a currency (Swiss franc) in the table will highlight the corresponding lollipop. Using Yi et al.’s taxonomy, we can label the interactions in both examples as connect: to “highlight associations and relationships between data items that are already represented” in different views (yi_toward_2007). However, if we focus on how the targets are changing, the two examples can be categorized differently. In Figure 3(a), the target (orange histograms) updates to show a subset of the data matching a specific month, hence it can be categorized as filter: “change the set of data items being presented based on some specific conditions” (yi_toward_2007); in Figure 3(b), on the other hand, the target (lollipop) is just a different visual representation of the same data item (Switzerland Franc), and it is differentiated from the other lollipops using fill color, hence matching the definition of select: “by making items of interest visually distinctive, users can easily keep track of them even in a large data set and/or with changes in representations” (yi_toward_2007).
These categories are from the same taxonomy, yet they do not focus on the same aspect of the interaction design: connect emphasizes the relationship between the mark clicked on and the target, without describing the effects on the target; whereas filter and select focus on what kind of change the target exhibits. The same issue applies to Heer and Shneiderman’s taxonomy (e.g., coordinate vs. filter).
5.2. A Taxonomy of Intent
Considering these two issues raised before, we decide to create our own list of categories for authoring intent and user intent by reinterpreting these existing taxonomies and analyzing examples. To do so, we need more precisely define authoring intent. When an author expresses her intent to author an interaction by connecting items, multiple aspects, derived from interactions between a user and a system, may need to be specified:
  • how the interaction is triggered (e.g., click, scroll, or UI control)
  • the type of target (e.g., mark, view) that exhibits visual change
  • the type of visual change (e.g., reorder, data updated) in target
  • whether the visual change triggered by interaction takes place within one view or across multiple views
Among these aspects, we argue the type of change in target is the most critical factor: it is the most immediately observable consequence of an interaction. We thus refine the definition of authoring intent as: what an author wants to achieve in terms of the desired type of change on visual objects by connecting items.
Our final taxonomy of authoring intent are new categories formulated based on the consolidation of user intents from existing taxonomies as well as new categories from our analysis. First of all, we collected a list of user intents related to how target object change (e.g., reconfigure, encode, filter). Yi et al.’s work do focus on such a perspective; hence, we mainly referred to the user intents from their taxonomy. We also made modifications based on the considerations discussed above. For example, we removed connect and explore, and added categories such as derive and annotate from Heer and Shneiderman’s taxonomy and Brehmer and Munzner’s typology. With the consolidated user intents, we grouped them by way of connect as authoring intents.
Here we present the categories: Authoring Intents (AI) and related user intents, and their definitions. We will elaborate on these categories with representative techniques and low-level semantic components in Section 8.
AI1: Connects visual objects and changes in visual properties. link user inputs to related system components to evoke changes in the visual aspects of target objects, thereby realizing these user intents:
  • Select: visually differentiate objects from their surroundings
  • Annotate: show auxiliary visual objects
  • Reconfigure: change the spatial arrangement of existing objects
AI2: Connects visual objects and changes in visible area. link user inputs to related system components to evoke changes in the range and viewpoint of visible area, thereby realizing this user intent:
  • Steer: direct camera view towards certain objects
AI3: Connects visual objects and changes in visual mapping. link user inputs to related system components to evoke changes in pairs of targets and corresponding data items, thereby realizing these user intents:
  • Encode: change the visual representation of existing data items
  • Enter/Exit: add or delete visual objects representing data items or fields to the visualization
AI4: Connects visual objects and changes in data. link user inputs to related system components to evoke changes in the underlying data engaged in targets’ visual updates, thereby realizing these user intents:
  • Filter: show a subset of the data items based on certain criteria
  • Abstract/Elaborate: change the levels of abstraction, which determines the data items to be visualized
  • Derive: compute and show new field values from data items attached to existing objects
In some cases regarding user intents, multiple interpretations are possible. For example, when filtering a stacked area chart, the spatial arrangement of the remaining area marks will change. We categorize such cases as filter instead of reconfigure, because filtering is the driving force behind the observed change, and the layout mechanism is unchanged.
6. Technique Categories
We consolidate techniques from existing taxonomies and incorporate additional techniques from our example analysis. Our primary motivation is to reference taxonomies that focus on target change-centric perspective when defining intents. Thus, we also began the amalgamation process using Yi’s and Heer and Shneiderman’s taxonomies (yi_toward_2007; heer_interactive_2012). Our initial step was to gather techniques associated with user intents from each paper. The Table 1 shows the collection. In consolidating these techniques, we further segment them into more granular levels. For example, while both papers define Select as a user intent, they do not delve into the detailed techniques. To address this, we include the related techniques with the user intent, Select (i.e., Point Select, Multi-Select, Range Select, Generalized Select, Linked Select, and Deselect), all informed by our example analysis. More techniques can be added as needed; thus, we define a group of primary techniques in Section 8, presenting a unified analysis.
Yi et al. (yi_toward_2007) Heer and Shneiderman (heer_interactive_2012)
User Intents Techniques User Intents Techniques
Select Select
Explore Pan, Direct Walk Navigate Pan, Scroll, Zoom
Reconfigure Sort, Align/Change Baseline, Change X/Y Fields in Chart, Change Attributes Assigned to Axis, Relocate an Item, Rotate Portion of the Visualization Sort Order, Reorder Data by Variables
Encode Semantic Zoom, Change Visual Properties, Change Chart Types Visualize Assign Data Variables to Visual Objects, Add an Item
Connect Highlight Connected Items, Reveal/Expand Related Items, Hide Unrelated Items Coordinate Brushing and Linking
Abstract/Elaborate Geometric Zoom, Tooltip Derive Nesting Variables
Organize Simplify Organization of Items, Switch Positions of Items, Resize a View
Record Undo and Redo Actions, Add Metadata, Bookmark
Annotate Textual Annotation, Draw a Circle/Arrow
Share Export Views or Data, Publish a Visualization
Table 1. User Intents and Techniques in Yi’s and Heer and Shneiderman’s taxonomies
7. Component Categories
In this section, we present the details on our analysis at the component level. The focus of our analysis at this level is: given a triggering mouse or keyboard event and a target, to attain the observed visual changes in the target, what visualization components are involved, and how are the components used or updated? We first review recent literature that examines the design space of static data visualizations (zhicheng_liu_atlas_2021; satyanarayan_critical_2019; liu_data_2018; ren_charticulator_2018; thompson_understanding_2020) and synthesize the components identified in these papers. Given the static components, we then used Vega-Lite’s interaction grammar to guide our top-down analysis of interactive behavior. In particular, we found the selection primitives to be powerful concepts to describe interaction behavior. Our analysis adapted and extends these concepts in two ways. First, we try to identify the components of interactivity with a focus on the input and output of each component. Second, the Vega-Lite does not provide an in-depth account of the role of UI controls, which are commonly used in real-world examples (26 out of 47 in our collection), and we introduced field and component references to address this issue. Below we introduce the component types uncovered from our analysis. We will elaborate on how the manipulation of these components relate to higher-level techniques and intents in Section 8.
7.1. Basic Components of a Static Representation
Visual Object. The following list specifies different types of visual objects in a visualization, going up the scene graph hierarchy:
  • mark: graphical elements such as shapes (rectangle, arc, line), symbols, images and texts;
  • glyph: a group of marks, for example, a lollipop chart (Figure 3(b)) uses glyphs consisting of one circle and a line;
  • collection: mark or glyph instances generated by data, for example, a simple bar chart contains one collection of bars, each bar represents a data item; collections can be nested, for example, a small multiple consists of a collection of mark collections;
  • axis and legend: reference objects that explain how to interpret a visualization;
  • annotation: auxiliary visual objects that provide contextual information or facilitate perceptual estimation;
  • scene: a self-contained view consisting of mark/glyph collections, axes, and legend. In a multiple coordinated view design, for example, each view is a scene. Similarly, in a tabbed interface, clicking on each tab shows one scene at a time;
  • section: multiple scenes can be organized into a section, a typical practice in scrollytelling interfaces where a narrative covers multiple topics, each topic involving a set of scenes.
Visual Channel. A visual object has a number of visual channels such as position, size, and color that can be used to represent data values or customize visual styles.
Data. Visual objects often have attached data items, which are encoded by properties (e.g., position, size) of the visual objects
Encoding. A declarative specification of the mapping between a data field and a visual channel.
7.2. Observable Interaction Components
In addition to components for static representations, we identify three components related to interaction that are visually observable: event, hit object, and target.
Event. Interactions are triggered by events through an input device. A mouse event consists of the following types: click, move, drag, and each event has associated parameters such as the x and y coordinates.
Listener. Listeners monitor and respond to events. The most used listener is the visualization canvas, which can be implemented as an SVG element or an HTML canvas. UI controls such as buttons and dropdown menus are also common listeners.
Hit Object. Given a mouse event received by a canvas listener, most visualizations conduct a hit-test (foley_computer_1996) to determine if any visual object intersects with mouse cursor. The result of the test, if not null, is a hit object. A hit object is usually a mark or a glyph.
Target. Targets are visual objects that undergo observable visual changes in response to a triggering event. In many examples, the hit object is a subset of the targets (e.g., in Figure 2, given a collection of circles, hovering over one of circles makes it a hit object, and all the circles in the collection are the targets, whose color and opacity are updated to highlight the hit object). An interactive visualization can also have targets without a hit object present: for example, we can achieve the same effect in Figure 2 by choosing a station from a drop-down menu.
7.3. Internal Components
Our analysis has identified the following components related to the internal working mechanisms of an interactive visualization. In contrast to the observable interaction components, these components have no visual representations in the user interface.
Camera. A component providing a vantage point of the visual representation. A camera has a number of configurable attributes, such as focus point (center of camera), field of view (extent visible through the camera), zoom level, and rotation. In web-based 2D visualizations, the camera is often implemented as the SVG view box (mdn_web_docs_viewbox_nodate) that can pan and zoom; in 3D visualizations, perspective cameras are used widely (sciencedirect_engineering_topics_perspective_nodate).
State Variables. An interactive visualization often requires state variables to configure and manage its state. Events update the values of these variables, which in turn trigger visual changes in target objects. Below is a list of exemplar state variables.
  • Predicate An expression consisting of a variable, an operator, and an operand value. The variable is usually a data field. For example, we can define a predicate country ==USA” where the operand is a single data value and the operator is equal; in another predicate, country $\in$ [“USA” “Canada”, “Mexico”], the operand is a list of values, and the operator is in. In some cases, the variable can be a visual property, e.g., x_position $\in$ [25, 50] if we want to do a collision check. A predicate can be evaluated to either TRUE or FALSE. They can thus be used to either retrieve data items, or test if a target satisfies a condition.
  • Field Reference Sometimes a visualization needs to keep track of the data field used to manipulate target objects. For example, in a bar chart where the bars can be sorted by different data fields, the visualization needs a sortBy variable to keep track of the data field determining the order.
  • Component Reference Direct references to visualization components are useful for navigational and presentation purposes. For example, in a scrollytelling interface where multiple scenes are included in the same page and shown at a predefined cadence, an state variable currentScene is necessary to decide which scene to display, and this variable can be bound to UI navigation controls like breadcrumbs as well as page view port.
Target Evaluator. A target evaluator is a function that takes a list of visual objects as input, and outputs numeric or Boolean values for each object based on the visual object’s data or visual properties. A target evaluator often involves one or more state variables. Below is a list of exemplar target evaluators:
  • A predicate evaluator tests if visual objects satisfy a given predicate. For example, given a bar chart where each bar represents a country and its height represents the population of each country, we can define an evaluator that tests each bar against the predicate country ==USA”, and the function will return a list of Boolean values, with TRUE for the bar representing USA only.
  • An order evaluator sorts a list of visual objects based on a sorting criteria, which can be expressed as a field reference, and returns a list of object indices.
  • A distance evaluator computes the pairwise similarity or distance between two objects. In a fisheye view, for example, given a hit object, we can define an evaluator that returns the index distance between each target and the hit object.
Evaluation Scale. A special type of scale specifying how to update target visual objects based on target evaluation results. Unlike a regular scale which takes data values as input, an evaluation scale takes results from a target evaluator as input. In the bar chart example above involving a predicate evaluator, the bar satisfying the predicate can be assigned a fill color different from the other bars.
8. Intents, Techniques, and Components: A Unified Account
In this section, we present a unified account of representative interaction techniques in our example collection, grouped by authoring intent and user intent in each subsection, with the low-level operation mechanisms described using static and interaction components. For each technique, we indicate whether it is primarily for a single view (S) or multiple views (M). We also use colored icons to highlight the components involved in each technique. When a component is in square brackets (e.g., [[Uncaptioned image]]), it means the component is optional. When multiple components are in parenthesis and separated by a vertical bar (e.g., ([Uncaptioned image][Uncaptioned image]) ), it means only one of the components is needed. It should be noted that the techniques presented here are not meant to be exhaustive.
8.1. Connects visual objects and changes in visual properties
8.1.1. Select: Visually Differentiate Objects from Surroundings
Point Select (S) & Multi-Select (S)
([Uncaptioned image][Uncaptioned image]) [Uncaptioned image] [Uncaptioned image]
In point selection, a visual object is marked and visually differentiated from its surroundings. For example, in Figure LABEL:fig:pointselect (alexander_bank_nodate), clicking on a glyph representing the date 12/21/2017 changes its stroke color and width, and makes the other glyphs opaque. To support point select, evaluate every target visual object to determine if it is the hit object; alternatively, create a predicate (e.g., data == 12/21/2017), and evaluate the targets against the predicate. The evaluation result is then used to encode the visual channels. Multi-select works in a similar way, except that the predicate operand is a list of data values, and the operator is in. For example, in Figure LABEL:fig:multiselect (parlapiano_how_2014), six selected polylines representing six industries are highlighted in green, while the other polylines turn gray and opaque. Point select and multi-select can be triggered by mouse events such as clicking (Figure LABEL:fig:pointselect) and scrolling (Figure LABEL:fig:multiselect), as well as UI events such as a button click.
Range Select (S)
([Uncaptioned image][Uncaptioned image]) [Uncaptioned image] [Uncaptioned image]
To range select visual objects along the x or y dimension (e.g., Figure 4(a)) (vega-lite_seattle_nodate), either (1) evaluate every target visual object to determine if it intersects with pixel range defined by the mouse event (click+drag); or (2) identify the data field encoding the position (e.g., date in Figure 4(a)), convert the pixel range into a predicate (e.g., date $\in$ [3/20, 8/6]), and evaluate if each target’s date value satisfies the predicate. The evaluation result is used to encode the visual channel values, where targets failing the evaluation turn gray.
Generalized Select (S, M)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Generalized selection (heer_generalized_2008) expands the hit object to a broader set of visual objects. For example, hovering over an arc representing a webpage click selects the arc as well as arcs representing the previous clicks in a sunburst chart (Figure 4(b)) (kerry_sequences_nodate). To support generalized selection, create a predicate based on the hit object (e.g., clickID $\in$ [sequence from root to “c5”]), then evaluate each target against the predicate, finally, encode the visual channel (i.e., opacity) values according to the evaluation result.
(a) Select a date range
(b) Select a tree branch
Figure 4. Range Select and Generalized Select
Linked Select (M)
([Uncaptioned image][Uncaptioned image]) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
To enable linked selection (e.g., Figure 3b) across multiple scenes, create a predicate based on either the hit object (if present) or the mouse coordinates, evaluate the target visual objects against the predicate, and encode the visual channel values according to the evaluation result.
Deselect (S, M)
[[Uncaptioned image]] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Two forms of deselection exist: users can click a selected object to remove it from selection, or they can click an empty spot in the background to clear selection. Both forms can be implemented as updating an existing predicate. For the former, remove the hit object ID from the predicate value operand; for the latter, set the predicate value operand to null. The evaluator is then invoked and the targets’ visual channels are encoded according to the evaluation results.
8.1.2. Annotate: Show Auxiliary Visual Objects
Show/Hide Reference Lines (S)
([Uncaptioned image][Uncaptioned image]) [[Uncaptioned image]]
A reference line is often shown to indicate the predicate value used for selection, or to facilitate visual estimation of data values. For example, in Figure 5(a) (wolf_alpha_nodate), hovering on a set of related charts sharing the same x-axis shows a vertical line to help users keep track of the location of their action and the currently selected date. In another example (Figure 5(b)) (scatterplot_with_voronoi_life_nodate), hovering over a circle shows two orthogonal orange lines, intended to help users read the x and y values more accurately. In our analysis, we assume that these reference lines have been created, and interaction primarily changes their visual channels such as visibility and position. To enable updating reference object, different components are involved depending on the visual channels and chart design. For example, in Figure 5(a), the vertical line’s position is based on mouse x-coordinate, and its visibility depends on whether a valid date predicate operand is found based on mouse x-coordinate; in Figure 5(b), the lines’ positions are based on mouse coordinates, and their visibility depends on whether a hit object is found.
(a) Indicate selected date
(b) Facilitate value reading
Figure 5. Annotate: Show Reference Lines
Show/Hide Tooltip Container (S)
([Uncaptioned image][Uncaptioned image]) [[Uncaptioned image]]
We define tooltips broadly to include visual objects that appear when hovering over a glyph or scene to provide contextual details about data items of interest. Tooltips thus include both the balloon element with text content “South Afria” in Figure 5(b) as well as the colored text labels in Figure 5(a) showing percentage values. To show a tooltip, typically a hit object is involved, and the position of the tooltip depends either on the mouse coordinates or the position of the hit object (Figure 5(b)). A tooltip can also be shown without any hit object, if a predicate is present (Figure 5(a)). Note that for the annotate intent, we only focus on the display of auxiliary visual objects, not their contents. Updating content is addressed in the Details on Demand technique under the filter intent category, which we will discuss in the next subsection.
8.1.3. Reconfigure: Change the Spatial Arrangement of Existing Objects
Reposition (S, M)
([Uncaptioned image][Uncaptioned image]) In scrollytelling visualizations, it is essential to control the pace of the narrative and dynamically update the positions of visual objects. In many cases, the new position is determined by the distance scrolled by mouse; in other cases, a target’s new position can be predefined and stored in a state variable. The second case can also be done by mouse click on a UI such as breadcrumb.
Sort (S)
[Uncaptioned image] [Uncaptioned image]
Sorting changes the order of visual objects based on field values. To enable sorting, a field reference orderBy needs to be defined, and its value can be updated by UI controls such as a dropdown menu. An order evaluator then computes the new list of object indices.
Organize Views (M)
[Uncaptioned image] [Uncaptioned image] Organize views is a technique to change the positions and sizes of scenes in a composite visualization. For example, in a multiple coordinated view design, by clicking on a view and dragging them to a new position, the view layout can be adjusted based on a layout evaluator embedded in the visualization system.
8.2. Connects visual objects and changes in visible area
8.2.1. Steer: Direct Camera View
towards Certain Objects
Geometric Zoom (S)
[Uncaptioned image] [Uncaptioned image]
Geometric zoom updates the visible area of a scene by reconfiguring camera focus point and zoom level based on mouse events. Figure 6 (bostock_zoomable_nodate) shows a zoomable bar chart. Users double click anywhere in the chart or pinch through mouse wheel to perform zooming. To enable geometric zoom, the coordinates of the input mouse event typically become the new focus point. The zoom level is usually updated by fixed increments based on the amount of pinching or the number of double clicking.
Figure 6. Zoom in and out on a bar chart
Pan (S)
[Uncaptioned image] [Uncaptioned image]
Unlike geometric zoom which can update both a camera’s focus point and zoom level, panning only changes the focus point based on mouse coordinates. Clicking & Dragging is the most common mouse event to perform panning (e.g., Figure 6) Both geometric zooming and panning take place at the scene level without a hit visual object, and everything inside the scene becomes a target.
Toggle between Different Views (M)
([Uncaptioned image][Uncaptioned image]) In tab-based user interfaces, it is common to display a view at a time by clicking on the UI control. For example, in GOV — DNA, by clicking on one of the four buttons at the top (e.g., Figure 8(a)), the currently visible view is updated. The views are usually predefined, each containing unique definitions of components, such as marks, encodings, and even predicates (to control which data items are visualized). Depending on how the views are defined, this technique can be realized by updating either a state variable or the camera. In Figure 8(a), the views are significantly different, and it is cleaner to treat them as independent scene definitions. Toggling between them thus requires the definition of a state variable currentScene. In examples such as the Star Mapper (interactive_esa_nodate), the toggling is more easily described by updating the camera’s position and focus point in a three-dimensional scene.
Navigate to a Scene or Section (M)
[Uncaptioned image] [[Uncaptioned image]]
Similar to Toggle between Different Views, this technique changes the current scene or section; however, this technique is primarily used in narrative visualizations. For example, based on the mouse coordinates scrolled (e.g., Figure 7) or the UI control clicked on (e.g., breadcrumb in Figure LABEL:fig:multiselect), the component reference currentScene updates, changing the currently visible view.
Figure 7. Navigating to the next section in a narrative visualization
(a) Change field in size encoding
(b) Change from multi-line chart to stacked area chart
Figure 8. Encode: Change Field and Chart Type
8.3. Connects visual objects and changes in visual mapping
8.3.1. Encode: Change the Visual Representation of Existing Data Items
Change Field in Encoding (S)
[Uncaptioned image] [Uncaptioned image]
In Figure 8(a), users can click on the left/right arrow buttons to choose a data field to encode the size of circles. Assuming that an encoder has already been created to map data values to a given visual channel properties, to enable changing field in encoding, we do not need to update the target’s data, we simply update the data field in the encoding component definition.
Change Chart Type (S)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
In Figure 8(b) (laura_namegrapher_nodate), clicking on the buttons (“Total” or “Compare”) toggles the visualization between a multi-line graph and a stacked area chart, showing the same data items. Multiple aspects of visual encoding are changed: mark type, visual channels and layout. To implement changing chart type, define state variables referencing the chart type definition and associated encodings, and create encoders that handle the mapping between data values to the corresponding marks and channels.
8.3.2. Enter/Exit
Click to Add Data Points (S)
In our example collection, we only encountered one case that introduces new data items through interaction. In a D3.js demo (mouseEvent_Will), clicking the background will add data points which are shown as new circle marks at the clicked locations. This example shows how interaction can modify the data being visualized.
8.4. Connects visual objects and changes in data
8.4.1. Filter
Dynamic Queries (S, M)
[Uncaptioned image] ([Uncaptioned image] [Uncaptioned image][Uncaptioned image])
Dynamic Queries (ahlberg_dynamic_1992; shneiderman_dynamic_1994) allow users to control which data items should be visualized through interaction with user interface elements such as sliders, checkboxes, and manipulable widgets (hochheiser_dynamic_2004; wattenberg_sketching_2001) that embody filtering criteria. To support dynamic queries, define a predicate for each querying control, and update predicate operands based on user interaction. The target visual objects can then be updated in two ways. First, we can evaluate the targets against a predicate evaluator, and encode the targets’ visibility accordingly, but this approach will not work for visualizations that require re-computation of layouts (e.g., stacked area chart in Figure 8(b)). Second, we can treat the scene as the target and update its data based on the predicate. The scene can then be updated with the same encoding mechanisms.
Details-on-Demand (S)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 2 is a canonical example of showing details on demand: when hovering over a mark, the tooltip content updates to show detailed information about the item of interest. We assume that the text elements in the tooltip and the mapping between the text elements and corresponding data fields have been created. Updating the content thus requires retrieving the relevant data items, which can be categorized as filter. To support details-on-demand, obtain a predicate from the hit object, and set the tooltip’s data as the data items satisfying the predicate. In our analysis, handling the visibility and position of the tooltip is a separate operation, which falls under the annotate category, as discussed in Section 8.1.2.
Cross-filter (M)
[[Uncaptioned image]] [Uncaptioned image] ([Uncaptioned image] [Uncaptioned image][Uncaptioned image])
Cross-filter occurs when user interaction in one scene filters the data represented by visual objects in other scenes (Figure 3(a)). The implementation of cross-filter is similar to that of dynamic queries: create a predicate based on input event, then either evaluate the target objects against the predicate and encode the evaluation results accordingly, or change the target scene’s data.
8.4.2. Abstract/Elaborate: Change the Levels of Abstraction
Move Up/Down in a Hierarchy
[Uncaptioned image] [Uncaptioned image]
Given a hierarchical dataset, users can choose whether to display the entire structure (including all the leaf nodes and nodes at the intermediate levels), or only nodes above a certain level. For example, DendroMap (bertucci_dendromap_2022) displays hierarchical clusters of image datasets. Users can increase or decrease the number of visible clusters (Figure 9) through a slider. The interaction sets the level of abstraction in the underlying hierarchy. To enable moving up and down a hierarchy, create a state variable recording the desired level of abstraction, and update the target data whenever the variable value changes.
Figure 9. Change the level of abstraction in a hierarchy
Drill Down/Roll Up
[Uncaptioned image] [Uncaptioned image]
Levels of abstraction can be changed on multivariate dataset too. In Polestar (moritz_polestar_nodate), Polaris (stolte_polaris_2002) and Tableau (tableau_software_tableau_nodate), users can drill down data by dragging additional dimensions to the row or column shelf, which will partition existing data displays into a finer grained aggregation. To enable drill-down and roll-up, create state variables recording the current dimensions and measure, and update the target data (based on pre-computed data cubes or real-time querying) whenever the variable values change.
8.4.3. Derive: Compute and Show New Field Values from Data Items Attached to Existing Objects
Recompute Field with New Baseline (S)
[Uncaptioned image] [Uncaptioned image]
Many visualizations show numeric measures computed against a baseline. Figure 3(b) shows the temporal trend of a currency’s (e.g., Swiss franc) Big Mac Index against a base currency (e.g., US dollar). Choosing a different baseline will generate a new set of index values that are encoded as the height of the lollipops. To enable re-computation against a dynamic baseline, define a state variable representing the baseline, and change the target scene’s data.
Figure 10. Change aggregator function to compute values
Change Aggregator (S)
[Uncaptioned image] [Uncaptioned image]
In visualizations where a mark represents multiple data items, it is possible to change how the values for a field are aggregated. For example, Polestar (moritz_polestar_nodate), Polaris (stolte_polaris_2002) and Tableau (tableau_software_tableau_nodate) allow users to choose from a list of aggregation functions (Figure 10). To support changing aggregator, define a state variable representing the aggregator, and update the target scene’s data whenever the aggregator is changed.
User Intent Technique Event Listener Hit Object Target Internal Components
Add Magnet enter add mark choose field dropdown menu none new magnet target data
Move Magnet reconfigure reposition click + drag canvas magnet all dust particles component references, target evaluator
Table 2. Dust and Magnet
User Intent Technique Event Listener Hit Object Target Internal Components
Add Matrix enter add collection choose set set list none new matrix target data
Move Matrix select point select click + drag canvas matrix all matrices predicate, target evaluator, evaluation scale
Compare Matrices abstract move up a hierarchy drag end canvas matrix all matrices target data
Change Operator derive change aggregator click canvas operator mark matrix state variable (logical operator), target data
Table 3. OnSet
8.5. Multiple User Intents
We also encountered many interaction designs demonstrating more than one intent. We present two such examples below.
Semantic Zoom: Steer + Abstract/Elaborate (S)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Unlike geometric zoom, semantic zoom has dual intents: steer and abstract/elaborate. When zooming in the heatmap in Figure 11(a) (heavyai_1990_nodate), not only the camera zoom level changes, the bin size of the heatmap also updates to show more granular information.
Direct Walk: Select + Steer (M)
([Uncaptioned image][Uncaptioned image]) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Direct walk “smoothly move[s] the viewing focus from one position in information structure to another” (yi_toward_2007). In Figure 11(b) (kerr_d3_nodate), clicking on a table row both selects a county map polygon and performs a geometric zoom with the county as the focus point.
(a) Semantic Zoom
(b) Direct Walk
Figure 11. Multiple Intents: Semantic Zoom and Direct Walk
9. Discussion: Descriptive, Evaluative, and Generative Powers and Completeness
We evaluate the descriptive, evaluative, and generative powers (beaudouin-lafon_designing_2004) of our framework, and discuss the completeness of the taxonomies.
9.1. Descriptive Power
To understand how well our framework can describe complex and bespoke interaction designs, we choose Dust and Magnet (DnM) (soo_yi_dust_2005) and OnSet (sadana_onset_2014) as two case studies, based on the following considerations: (1) they use customized static visual representations instead of standard chart types, and (2) the interactive behaviors deviate from the representative techniques included in our analysis.
Dust and Magnet. DnM allows users to interactively explore a multivariate dataset. Users can add one or more fields as magnets. All the data items are shown as iron particles. When moving a magnet representing a field, the dust particles are attracted to the magnet, where the particles with high values for the given field will move closer to the magnet. Table 3 analyzes two primary actions in DnM using our framework. For adding a magnet, The authoring intent is to link the user input to the relevant system components, thereby generating a new magnet. The user intent is to introduce a new field. Choosing an option from the dropdown menu adds a data field, which is visualized as a magnet. For moving a magnet, the authoring intent is to link the user input to the relevant system components, thereby changing positions of the magnet. The user intent is to reconfigure. Consequently, we need a target evaluator that takes all the data items as input, and computes the new positions for each item. To perform this computation, the target evaluator keeps track of all the references to the existing magnets to read their positions and fields.
OnSet. Developed to visualize sets, OnSet uses matrices to represent a single set, logical operations on multiple sets, or nested logical operations on multiple sets. Each data item is assigned a fixed position in the matrices, and its presence or absence in a set is shown using color. The appearance of a matrix varies accordingly to the data it represents (i.e., a single set or composite sets). Table 3 analyzes four primary actions supported in Onset, encompassing three authoring intents and four user intents. Both adding a matrix and comparing matrices modify the target data: the former adds a new set (a tree with only one node) which is visualized as a matrix; the latter constructs a tree where the leaf nodes are individual sets and their ancestors are results of applying logical operators on the children sets. Note that OnSet has defined encodings and scales for determining the visual representations of a tree depending on its structure and content, and these rules remain independent from the interactions. When moving a matrix (the hit object), a predicate is used by a target evaluator to determine if the bounding box of each matrix in the visualization intersects with the hit object, and an evaluation scale assigns visual properties to each matrix based on the evaluation result. Finally, to support toggling between “AND” and “OR” operators on a composite matrix, the program needs to keep a state variable for the operator and update the matrix’s data whenever the operator changes.
9.2. Evaluative Power
Our framework provides a theoretical lens and vocabulary to evaluate existing and new languages and tools for composing interactive visualizations. The evaluation can be conducted across three aspects: levels of abstraction, expressiveness, flexibility.
Levels of abstraction. The first aspect, levels of abstraction is to examine where a system primarily operates at. Each system encompasses different levels of abstraction, which defines the range of interaction authoring in the system. For example, Interactive+ is to augment an existing static web-based visualization with interactivity. This tool operates primarily at the component level: given a visualization realized in the HTML Document Object Model (whatwg_dom_nodate), Interaction+ extracts component-level information including the visual objects and their attributes as well as the distribution of visual objects over each visual attribute. Interaction+’s user interface does not provide explicit abstractions of concepts at the technique and user intent level. As an example of a system that covers multiple abstraction levels, consider Lyra 2, which grants access to two levels of abstraction. At the component level, author’s demonstration is primarily in the form of choosing visual marks in a static visualization generated by Vega or Vega-Lite, and constructing query objects that are translated to predicates. An interaction inspector also provides component-level information about recommended design (e.g., mouse event, visual attributes); at the technique level, the suggested interaction techniques are shown as selectable widgets with visual thumbnails in the inspector. Lyra 2 does not provide explicit visual representations or interaction mechanisms at the user intent level.
Expressiveness. The second aspect, expressiveness, evaluates the range of interactivity against the categories at the identified levels of abstraction. For instance, Interaction+ is tool-agnostic, allowing users to add interactions to a static web-based visualization without accessing the underlying data. Consequently, this tool supports some of the techniques such as generalized selection without a predicate component involving data updates. It does this by defining filtering criteria based on visual attributes alone. The exclusion of data engagement restricts expressiveness by not supporting some of the data-related techniques (e.g., Click to add data point, Change field in encoding, Recompute Field with New Baseline and more). Even for techniques such as generalized selection, if a data field is not visually encoded, it would not be possible to select based on the field in Interaction+. With Vega and Vega-Lite as the underlying grammars for static visualizations, Lyra 2 has access to a much more complete set of components than Interaction+ (e.g., encodings and data). It is thus more expressive, supporting techniques such as generalized selection involving data projection. However, some of richer techniques are unavailable (e.g., Navigate to Previous/Next Scene or Section, Organize view, and Reposition). As interactive visualizations have become highly expressive, they demand advanced techniques, such as scrolly-telling and drag-and-drop dashboards. Thus, with Lyra 2, how to compose such interactions remains unclear.
Flexibility. The third aspect, flexibility, refers to the extent to which tools or languages can be easily adjusted or expanded to accommodate users’ specific requirements. Take the example of Interaction+ as we previously discussed regarding its limitation in terms of expressiveness, the tool also exhibits inflexibility. This is primarily due to the restricted techniques stemming from omitting data when authoring interactions, the unclear definitions of their six categories of user intents, and the absence of UI abstractions for techniques. Such limitations make the tool difficult to flexibly operate across multi-levels of abstraction. Overall, Interaction+ is primarily created for use cases where interaction centers around visual objects and attributes, and it may not be sufficient to provide support for either fine-grained interaction design activities involving data and state variables, or concise specification of high-level authoring intents and techniques. Meanwhile, Lyra 2 exhibits a considerable amount of flexibility. By providing visual widgets representing the suggested interaction design, authors can think and interact with the concepts at the technique level. They also can easily inspect and change lower-level components such as mouse events and visual channels. Users cannot specify interaction design at the intent level since the interface provides no abstractions at that level.
In this section, we showcase the evaluative power of our framework by comparing two examples with different levels of abstraction, expressiveness, and flexibility. Our framework occupies a middle ground between template-based tools (e.g., Tableau) and low-level visualization libraries (e.g., D3.js). To encompass all levels of abstraction represented in our framework, we selected two non-template-based authoring systems that incorporate multiple techniques and operate at a component level of abstraction. Based on our understanding, there are limited systems that fulfill these criteria. Therefore, we plan to develop an interaction authoring grammar to showcase our framework’s evaluative power further.
9.3. Generative Power
Our analysis identifies components associated with each technique, which can inform the implementation of interactivity in visualization. It also has the potential to help create new languages and systems for interaction authoring. We demonstrate the generative power of our framework by proposing new ideas and guidelines.
First of all, both Interaction+ and Lyra 2 focus on creating interactivity for marks, and mostly overlook the other types of visual objects such as collection and scene. By including these visual objects in the scope, techniques such as OnSet, scrollytelling and dynamic toggling are possible. Second, it is important that an expressive interaction authoring language or tool needs to understand and provide internal abstractions for various representation and interaction components. It would also be helpful to expose these components in the interface through manipulable visual representations by adopting, for example, a visual programming paradigm (kuhail_characterizing_2021). Finally, a powerful authoring tool can allow users to start at any of the three levels of abstraction, depending on their expertise and tasks, and intelligently provide assistance through appropriate abstractions. The introduction mentioned one example where an intent can be translated into techniques and low-level implementations. We can imagine a different scenario where an author has composed an interaction technique using certain components, and wants to see how the implementation can be generalized to other techniques and even intents.
9.4. Completeness
When selecting real-world examples, we sought to increase the diversity in terms of chart type and source (Section 4). We do not claim, however, that the categories at each level of our framework are comprehensive and complete, capturing all possible interaction designs. First of all, we can only include representative techniques for each intent category. Many interaction designs, including the two in Section 9.1, bear multiple intents. The list of techniques is thus certainly not exhaustive. For the intent and component categories, it is not our primary goal to ensure the taxonomies are complete. As demonstrated in scientific fields like biology where classifications of life forms are constantly updated, we anticipate that our taxonomies will evolve with new entries and revisions as novel interaction designs are created.
10. Conclusion
We present a unified framework of the task space of interaction authoring in data visualization at three levels: authoring intent, technique, and component. We identify conceptual categories at each level, and discuss how these categories can be linked for a holistic understanding of interaction design. This analysis represents a step towards a systematic theoretical framework for interactive visualization design.
Acknowledgements.
This research was conducted as part of the requirements for Hyemi Song’s Master’s degree in Computer Science at the University of Maryland, College Park, from Spring 2022 to Summer 2023. This paper was peer-reviewed and invited for fast-track submission to the Computer Graphics Forum (CGF) journal via EuroVis 2023.
Contact
For correspondence, please contact the first author (hsong02@cs.umd.edu) at the University of Maryland College Park.